# Luca Eyring1,2,3,  Shyamgopal Karthik2,3,4,11footnotemark: 1  Karsten Roth2,3,4 Alexey Dosovitskiy5Zeynep Akata1,2,3

### v1

@@ 3 Related Work
Initial Noise Optimization.
The initial noise optimization framework was first introduced in DOODL [81] for improved guidance in Text-to-Image models. Subsequently, it was leveraged by Karunratanakul et al. [38] for 3D universal motion priors, for rare-concept, generation in text-to-image models [70, 69, 54], and by D-Flow [5] for solving inverse problems in various settings. While these methods mainly focus on controlling the generated sample for specific applications, our proposed method is designed to generally improve Text-to-Image models without the need for additional techniques to mitigate exploding or vanishing gradients on the optimization process. Most related to our work is DOODL [81], which also proposes to improve the textual alignment of text-to-image models using a CLIP-score-based criterion function, which we similarly employ in our method. These existing methods, however, take 10 (DOODL) to 40 (D-Flow) minutes to generate a single image due to their application on time-dependent generative models with a large number of denoising steps. To mitigate this, Samuel et al. [70] propose a bootstrap-based method to increase the efficiency of generating a batch of images. However, this method is limited to settings where the goal is to generate samples including a concept jointly represented by a set of input images.
Reward Optimization for Text-to-Image Models.
Reward models [89, 41, 88, 87, 42] were first introduced to mimic human preferences given an input prompt and generated images. There have been several attempts at incorporating these signals to enhance text-to-image generation. One notable direction is the idea of using reinforcement learning based algorithms to fine-tune text-to-image models to better align with these rewards either with an explicit reward model [8, 21, 10, 16, 28, 95] or by bypassing it entirely with Direct Preference Optimization [64, 90, 45, 82]. However, this can be expensive, requiring thousands of queries to generalize, and therefore a lot of work has explored directly fine-tuning diffusion models [15, 46, 62] using differentiable rewards [89, 42, 88, 41, 35]. Additionally, there has also been works exploring the concept of using reward models to perform classifier-guidance [4, 30] as well as using rewards to distill diffusion models into fewer steps [43, 65]. Differently from these works, we focus on adapting a diffusion model during inference by purely optimizing the initial latent noise using a differentiable objective.
@@ 4 Experiments
Experimental Setup.
We extensively evaluate the effectiveness of our proposed method, ReNO, using four open-source one-step image generation models: SD-Turbo, SDXL-Turbo, PixArt-$\alpha$ DMD, and HyperSDXL. HyperSDXL generates images of size $1024\times 1024$ while the others generate $512\times 512$. To assess the performance across diverse scenarios, we consider three challenging tasks. First, we evaluate on T2I-CompBench [33], which comprises 6000 compositional prompts spanning six categories, using a VQA, object detection, and image-text matching scores. Second, we employ GenEval [26], consisting of 552 object-focused prompts, measuring the quality of the generated images using a pre-trained object detector. Finally, we utilize Parti-Prompts [92], a collection of more than 1600 complex prompts, and assess the generated images using both reward-based metrics and extensive human evaluation. Throughout all experiments, we optimize Equation (7) for 50 steps using gradient ascent with Nesterov momentum and gradient norm clipping for stability. Lastly, we select the image with the highest reward score from the optimization trajectory for evaluation.
@@ 4.1 Effect of Reward Models
Table 2: Quantitative Results on T2I-CompBench. ReNO combined with (1) PixArt-$\alpha$ DMD [12, 11, 91], (2) SD-Turbo [72], (3) SDXL-Turbo [72], HyperSD [65] demonstrates superior compositional generation ability in both attribute binding, object relationships, and complex compositions. The best value is bolded, and the second-best value is underlined. Multi-step results taken from [12, 20].
@@
Table 3: Quantitative Results on GenEval. ReNO combined with (1) PixArt-$\alpha$ DMD [12, 11, 91], (2) SD-Turbo [72], (3) SDXL-Turbo [72], HyperSDXL [65] improves results across all categories. The best value is bolded, and the second-best value is underlined. Multi-step results taken from [20].
@@ 4.2 Quantitative Results
Table 2 presents the quantitative results of ReNO on T2I-Compbench. Most notably, we observe that for both Pixart-$\alpha$ DMD and SD-Turbo, there are improvements of over 20% in the Color, Shape, and Texture Categories. For instance, on Color SD-Turbo improves from 55% to 78%, which is only slightly below DALL-E 3. Similar improvements can also be seen for SDXL-Turbo and HyperSDXL models where performance increases by over 10 percentage points in these categories. Even outside this, there are significant boosts in the Spatial, Non-Spatial, and Complex categories, highlighting both the efficacy of the noise optimization framework, as well as the utility of human preference models for improving T2I generation at inference. Similar trends can also be noticed for GenEval in Table 3, where applying our noise optimization framework helps improve the performance of various one-step diffusion models. For instance, SD-Turbo improves its mean score from 0.49 to 0.62. Notably, our strongest model, HyperSDXL + ReNO, comes very close to the proprietary DALL-E 3 and SD3, i.e., beating DALL-E 3 on 4/6 categories in GenEval. In both of these benchmarks, our noise optimization framework improves results for all the models in all the categories. It is also important to note that both T2I-Compbench and GenEval use a variety of methods unrelated to human preference rewards, such as VQA models and object detectors, to detect different objects in the generated images. We report further quantitative results including comparisons to other test-time-based methods in Appendix B. Additionally, these quantitative results are supported by the qualitative results reported in Figure 1 and Appendix A.
@@ 4.4 Computational Cost of ReNO
Limitations.
An interesting observation in our experiments is that despite using different image generation models of varying architectures and sizes, they converge to similar performance on both T2I-Compbench and GenEval. In addition to the limitations of the generative models, we hypothesize that this could be due to the limitations of the reward models themselves, given their limited compositional reasoning abilities [93]. Stronger reward models and preference data would be crucial in enhancing results further.
@@
Method sec/iter (total) VRAM #params Img size SD-Turbo 0.4s (20s) 15GB 860M $512\times 512$ SDXL-Turbo 0.6s (30s) 21GB 2.6B $512\times 512$ PixArt-$\alpha$ DMD 0.5s (25s) 25GB 600M $512\times 512$ HyperSDXL 1.0s (50s) 39GB 2.6B $1024\times 1024$
@@ 5 Conclusion
This work introduces ReNO, a test-time optimization strategy for enhancing text-to-image generation without any fine-tuning. Not only do we achieve the strongest results among all open-source models on T2I-Compbench and GenEval, but images from ReNO on a single-step SD-Turbo have over a 60% win rate against a 50-step SDXL model and is competitive with the 8B parameter SD3 model on user studies. We also demonstrate that ReNO outperforms SDXL even when restricted to the same computational budget, highlighting the benefits of ReNO for practical use cases. The performance gains from ReNO underscore the importance of developing even better and more robust reward models and, moreover, establish a valuable benchmark for assessing their effectiveness. Furthermore, the substantial impact of optimizing the initial noise distribution motivates further research into understanding, manipulating, and controlling this crucial aspect of generative models.
@@ Appendix A Further Qualitative Results
Here, we report further qualitative results. We separate them into $512\times 512$ and $1024\times 1024$ generated images. First, in Figure 7, we show examples of ReNO applied to SD-Turbo, SDXL-Turbo, and PixArt-$\alpha$ DMD. Then, in Figures 8 and 9, we report HyperSDXL + ReNO against competing methods. Broadly, we see that ReNO not only fixes the artifacts occurring in one-step models but also improves compositional understanding (color attribution, spatial reasoning) as well as the quality of generated faces and pushes current one-step models to be broadly on par with proprietary models in these settings.
@@ B.1 Comparison to Direct Preference Optimization
B.2 Compositional Text-to-Image Methods.
@@ Appendix D User Study
Appendix E Algorithm
@@
Appendix F Implementation Details
F.1 ReNO hyperparameters
@@ F.3 Rewards
F.3.1 Human Preference Score v2 (HPSv2)
@@ F.4.2 GenEval
F.4.3 LAION Aesthetic Score Predictor
@@
Appendix G Broader Impact

### v2

# Luca Eyring1,2,3,  Shyamgopal Karthik1,2,3,411footnotemark: 1  Karsten Roth2,3,4 Alexey Dosovitskiy5Zeynep Akata1,2,3

@@ 3 Related Work
Initial Noise Optimization. The initial noise optimization framework was first introduced in DOODL [91] for improved guidance in Text-to-Image models. Subsequently, it was leveraged by Karunratanakul et al. [43] for 3D universal motion priors, for rare-concept generation [80, 79, 61] and enhancing image quality [31, 89] in text-to-image models, music generation [63, 62], and by D-Flow [5] for solving inverse problems in various settings. While these methods mainly focus on controlling the generated sample for specific applications, our proposed method is designed to generally improve Text-to-Image models without the need for additional techniques to mitigate exploding or vanishing gradients on the optimization process. Most related to our work is DOODL [91], which also proposes to improve the textual alignment of text-to-image models using a CLIP-score-based criterion function, which we similarly employ in our method. These existing methods, however, take 10 (DOODL) to 40 (D-Flow) minutes to generate a single image due to their application on time-dependent generative models with a large number of denoising steps. To mitigate this, Samuel et al. [80] propose a bootstrap-based method to increase the efficiency of generating a batch of images. However, this method is limited to settings where the goal is to generate samples including a concept jointly represented by a set of input images.
Reward Optimization for Text-to-Image Models. Reward models [100, 46, 98, 97, 47] were first introduced to mimic human preferences given an input prompt and generated images. There have been several attempts at incorporating these signals to enhance text-to-image generation. One notable direction is the idea of using reinforcement learning based algorithms to fine-tune text-to-image models to better align with these rewards either with an explicit reward model [8, 23, 11, 18, 30, 109] or by bypassing it entirely with Direct Preference Optimization [74, 101, 50, 92]. However, this can be expensive, requiring thousands of queries to generalize, and therefore a lot of work has explored directly fine-tuning diffusion models [17, 51, 72] using differentiable rewards [100, 47, 98, 46, 39]. Additionally, there has also been works exploring the concept of using reward models to perform classifier-guidance [4, 34] as well as using rewards to distill diffusion models into fewer steps [48, 75]. Differently from these works, we focus on adapting a diffusion model during inference by purely optimizing the initial latent noise using a differentiable objective.
@@ 4 Experiments
Experimental Setup. We evaluate the effectiveness of our proposed method, ReNO, using four open-source one-step image generation models: SD-Turbo, SDXL-Turbo, PixArt-$\alpha$ DMD, FLUX-schnell and HyperSDXL. HyperSDXL generates images of size $1024\times 1024$ while the others generate $512\times 512$. To assess the performance across diverse scenarios, we consider three challenging tasks. First, we evaluate on T2I-CompBench [37], which comprises 6000 compositional prompts spanning six categories, using a VQA, object detection, and image-text matching scores. Second, we employ GenEval [28], consisting of 552 object-focused prompts, measuring the quality of the generated images using a pre-trained object detector. Finally, we utilize Parti-Prompts [103], a collection of more than 1600 complex prompts, and assess the generated images using both reward-based metrics and extensive human evaluation. Throughout all experiments, we optimize Equation (7) for 50 steps using gradient ascent with Nesterov momentum and gradient norm clipping for stability. Lastly, we select the image with the highest reward score from the optimization trajectory for evaluation.
@@ 4.1 Effect of Reward Models
Table 2: Quantitative Results on T2I-CompBench. ReNO combined with (1) PixArt-$\alpha$ DMD [13, 12, 102], (2) SD-Turbo [81], (3) SDXL-Turbo [81], (4) HyperSD [75] demonstrates superior compositional generation ability in both attribute binding, object relationships, and complex compositions. The best value is bolded, and the second-best value is underlined. Multi-step results taken from [13, 22].
@@
Table 3: Quantitative Results on GenEval. ReNO combined with (1) PixArt-$\alpha$ DMD [13, 12, 102], (2) SD-Turbo [81], (3) SDXL-Turbo [81], (4) HyperSDXL [75] improves results across all categories. The best value is bolded, and the second-best value is underlined. Multi-step results taken from [22].
@@
FLUX-schnell 0.64 0.98 0.80 0.64 0.78 0.18 0.43
FLUX-schnell + ReNO (Ours) 0.72 0.99 0.90 0.79 0.87 0.21 0.56
FLUX-dev 0.68 0.99 0.85 0.74 0.79 0.21 0.48
@@ 4.2 Quantitative Results
Table 2 presents the quantitative results of ReNO on T2I-Compbench. Most notably, we observe that for both Pixart-$\alpha$ DMD and SD-Turbo, there are improvements of over 20% in the Color, Shape, and Texture Categories. For instance, on Color SD-Turbo improves from 55% to 78%, which is only slightly below DALL-E 3. Similar improvements can also be seen for SDXL-Turbo and HyperSDXL models where performance increases by over 10 percentage points in these categories. Even outside this, there are significant boosts in the Spatial, Non-Spatial, and Complex categories, highlighting both the efficacy of the noise optimization framework, as well as the utility of human preference models for improving T2I generation at inference. Similar trends can also be noticed for GenEval in Table 3, where applying our noise optimization framework helps improve the performance of various one-step diffusion models. For instance, SD-Turbo improves its mean score from 0.49 to 0.62. Notably, our strongest model, HyperSDXL + ReNO, comes very close to the proprietary DALL-E 3 and SD3, i.e., beating DALL-E 3 on 4/6 categories in GenEval. In the case of FLUX-schnell, ReNO improves the performance (0.72) to even surpass that of the base FLUX-dev model (0.68). Most notably, this is the strongest open-source results reported on the GenEval benchmark. In both of these benchmarks, our noise optimization framework improves results for all the models in all the categories. It is also important to note that both T2I-Compbench and GenEval use a variety of methods unrelated to human preference rewards, such as VQA models and object detectors, to detect different objects in the generated images. We report further quantitative results including comparisons to other test-time-based methods in Appendix B. Additionally, these quantitative results are supported by the qualitative results reported in Figure 1 and Appendix A. Lastly, we report full details for the conducted FLUX-schnell experiments in Appendix E.8.
@@ 4.4 Computational Cost of ReNO
4.5 Effect of ReNO on the Diversity of Generated Images
Table 4: We measure the average LPIPS and DINO similarity scores over images generated for 50 different seeds for 100 prompts from Parti-Promtps.
LPIPS $\downarrow$ DINO $\downarrow$ SD-Turbo 0.382 ±$\pm 0.043$ 0.770 ±$\pm 0.101$ SD-Turbo + ReNO 0.246 ±$\pm 0.046$ 0.712 ±$\pm 0.132$ SD2.1 (50-step) 0.243 ±$\pm 0.049$ 0.623 ±$\pm 0.150$ SDXL-Turbo 0.391 ±$\pm 0.044$ 0.835 ±$\pm 0.073$ SDXL-Turbo + ReNO 0.291 ±$\pm 0.041$ 0.763 ±$\pm 0.116$ SDXL (50-step) 0.351 ±$\pm 0.042$ 0.700 ±$\pm 0.128$
To investigate the effect of noise optimization on output diversity, we evaluate images generated across 50 different random seeds for 110 prompts from Parti-Prompts. Specifically, we generate a batch of images and use LPIPS [106] and DINO [9, 64] scores to compute the average similarity of the generated batch, where a lower similarity score corresponds to higher diversity. As shown in Table 4, one-step models (SD-Turbo, SDXL-Turbo) exhibit lower diversity compared to their multi-step counterparts (SD2.1, SDXL), likely due to adversarial training. However, applying ReNO not only maintains but actually increases diversity. For both SD-Turbo and SDXL-Turbo, ReNO achieves diversity levels approaching their respective multi-step base models highlighting an unexpected benefit of noise optimization increasing diversity. Figure 10 illustrates these improvements qualitatively. We hypothesize that the reason for this increased diversity is that ReNO adds structure to the noise, thus optimizes it away from the zero mean of the noise distribution and creating more diverse noises compared to sampling from the standard Gaussian.
4.6 Comparison to Multi-Step Noise Optimization
We benchmark ReNO against DOODL [91], which performs noise optimization using the 50-step SD2.1 model. Due to DOODL’s computational demands, we evaluate on the first 50 prompts from T2I-CompBench. Despite using the same CLIPScore objective, ReNO achieves four times larger improvements in the optimized criterion while requiring 75% less GPU memory and running 100x faster, highlighting the effectiveness of our one-step approach. Moreover, ReNO’s multi-reward objective leads to substantially larger gains in attribute binding accuracy (21.0-28.9%) compared to solely using CLIPScore, reiterating the efficacy of ReNO’s optimization objective.
Table 5: Performance comparison of ReNO and DOODL over the first 50 prompts of each of the Attribute Binding categories in T2I-CompBench. We report scores from default T2I-Compbench evaluation using BLIP-VQA as well as the optimized CLIPScore before and after optimization.
[…]
4.7 Limitations
An interesting observation in our experiments is that despite using different image generation models of varying architectures and sizes, they broadly converge to similar performance on both T2I-Compbench and GenEval. In addition to the limitations of the generative models, we hypothesize that this could be due to the limitations of the reward models themselves, given their limited compositional reasoning abilities [104]. Stronger reward models [53, 107] and preference data [33, 108, 15, 42, 99] would be crucial in enhancing results further.
@@ 4.7 Limitations
Method sec/iter (total) VRAM #params Img size SD-Turbo 0.4s (20s) 15GB 860M $512\times 512$ SDXL-Turbo 0.6s (30s) 21GB 2.6B $512\times 512$ PixArt-$\alpha$ DMD 0.5s (25s) 25GB 600M $512\times 512$ FLUX-schnell 0.6s (30s) 50GB 12B $512\times 512$ HyperSDXL 1.0s (50s) 39GB 2.6B $1024\times 1024$
@@ 5 Conclusion
We introduce ReNO, a test-time optimization strategy for enhancing text-to-image generation without any fine-tuning. Not only do we achieve the strongest results among all open-source models on T2I-Compbench and GenEval, but images from ReNO on a single-step SD-Turbo have over a 60% win rate against a 50-step SDXL model and is competitive with the 8B parameter SD3 model on user studies. We also demonstrate that ReNO outperforms SDXL even when restricted to the same computational budget, highlighting the benefits of ReNO for practical use cases. The performance gains from ReNO underscore the importance of developing even better and more robust reward models and, moreover, establish a valuable benchmark for assessing their effectiveness. Furthermore, the substantial impact of optimizing the initial noise distribution motivates further research into understanding, manipulating, and controlling this crucial aspect of generative models.
@@ Acknowledgements
The Appendix is organized as follows:
  • Section A provide further qualitative results.
  • Section B presents further quantitative analysis.
  • Section C provides an analysis of the different employed reward models.
  • Section D describes the full details for our user study setup.
  • Section E outlines the implementation specifics, including the algorithm.
@@ Appendix A Further Qualitative Results
Here, we report further qualitative results. We separate them into $512\times 512$ and $1024\times 1024$ generated images. First, in Figure 7, we show examples of ReNO applied to SD-Turbo, SDXL-Turbo, and PixArt-$\alpha$ DMD. Then, in Figures 8 and 9, we report HyperSDXL + ReNO against competing methods. Broadly, we see that ReNO not only fixes the artifacts occurring in one-step models but also improves compositional understanding (color attribution, spatial reasoning) as well as the quality of generated faces and pushes current one-step models to be broadly on par with proprietary models in these settings. Lastly, we report qualitative results for the effect of ReNo on the diversity of generated images in Figure 10.
@@
Figure 10: Non-cherry-picked results for SD-Turbo with and without ReNO for two different prompts over the first 5 seeds. ReNO increases the diversity of generated images w.r.t. content and layout.
@@ B.1 Comparison to Direct Preference Optimization
Comparison to Multi-step Noise Optimization Methods.
In Section 4.6, we report quantitative comparison between the multi-step noise optimization method DOODL and ReNO. In Table 9, we additionally report more details on the difference in efficiency between DOODL and ReNO. Note that for the same objective and model family ReNO is $120$x faster compared to DOODL.
Table 9: Computational cost comparison of ReNO compared to DOODL.
Model sec/iter (total) T2I-CompBench VRAM #params
SD2.1 + DOODL (CLIP) 24s (20min) 83.33 A100 days 40GB 860M
SD-Turbo + ReNO (only CLIP) 0.2s (10s) 0.63 A100 days 10GB 860M
SD-Turbo + ReNO 0.4s (20s) 1.25 A100 days 15GB 860M
@@ Appendix D User Study
Appendix E Implementation Details
E.1 Algorithm
@@
E.2 ReNO hyperparameters
@@ E.4 Rewards
E.4.1 Human Preference Score v2 (HPSv2)
@@ E.5.2 GenEval
E.5.3 LAION Aesthetic Score Predictor
@@
E.6 Diversity Analysis
We generated images with 50 different seeds for 10 prompts from each of the 11 challenges of PartiPrompts, totaling 110 prompts. Then, for each prompt, we evaluate the diversity over the 50 seeds by computing the mean pairwise LPIPS [106] and DINO [9, 64] scores. The higher these two scores are, the less diverse the generated images across seeds. We report the mean and standard deviation across all prompts.
E.7 Comparison to DOODL
For our comparison to multi-step noise optimization DOODL, we use the first 50 prompts from each of the attribute binding categories of T2I-CompBench. We benchmark DOODL using the official codebase (https://github.com/salesforce/DOODL/blob/main/doodl.py), adapted to SD2.1 with 50 steps. We chose to focus on the SD2.1 model family because when running DOODL on SDXL, it exceeds 40GB of VRAM making it unfeasible for single GPU runs and thus inference.
E.8 FLUX-schnell results
We find that noises employed for FLUX-schnell with one step translate very well to FLUX-schnell wtih four steps. Thus, due to efficiency we apply ReNO to FLUX-schnell with one step and afterward feed in the optimal noise to the four step FLUX-schnell model to obtain our final generation. Due to VRAM constraints, we generate samples in $512\times 512$ including CPU-offloading such that FLUX-schnell + ReNO runs within 40GB of VRAM. The FLUX-dev results reported in Table 3 are taken from Liu et al. [55]. We report FLUX-schnell + ReNO results on the attribute binding categories of T2I-CompBench in Table 12 and a qualitative comparison in Figure 12.
Table 12: Comparison of $512\times 512$ FLUX-schnell with and without ReNO on the attribute binding categories of T2I-CompBench.
Attribute Binding
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
FLUX-schnell 0.69 0.53 0.67
FLUX-schnell + ReNO 0.80 0.60 0.75
Figure 12: Non-cherry-picked results ($seed=0$) for FLUX-schnell with and without ReNO compared to FLUX-dev.
Appendix F Broader Impact